Nine boys and twelve girls have signed up for a trip. Only six students will be selected to go on the trip. Determine the probability that only boys will end up on the trip.
Thank you in advance for any help:)
I think the answer is : 0.08%

Answer by math_helper(2461)


9 boys + 12 girls = 21 students total
P(out of 6 students selected at random, all are boys) =
(9/21)(8/20)(7/19)(6/18)(5/17)(4/16) = 60480 / 39070080
= 1/646 or approximately 0.00155 (0.155%)